im dealing wit slopes. i have been doing problems like solve for y when slope is 4 passing through (6,7) now i have a problm that says my slope is 4/5 passing thru the origin. wat is the origin?
Answer by nerdybill(7384)

The "origin" is at (0,0) the center of the x-y coordinate axis.
.
So, given:
slope as 4/5
and crossing (0,0)
.
The "slope" intercept form of a line is:
y = mx + b
where
m is the slope
b is the y-intercept
.
If your line crosses (0,0) the y-intercept is at 0.
Plug the given info into:
y = mx + b
y = (4/5)x + 0
y = (4/5)x (this is the formula of the line in "slope-intercept" form)